# Superfast Low Rank Approximation

### v2

@@ 1 Introduction
1.6. Organization of our paper. In the next subsection we recall some definitions and auxiliary results. In Part I we cover computation of CUR LRA. We recall some background material in the next two sections. In Sec. 4 we estimate output errors of CUR LRA, proving our Main Result I. In Sec. 5 we cover numerical experiments for the algorithms of Part I. In Part II (Secs. 69) we study random sketching algorithms for LRA. In Secs. 6 and 7 we recall and slightly extend some known ones and recall their deterministic output error bounds, respectively. In Sec. 8 we prove error bounds for our dual LRA algorithms – our Main Result 2. In Sec. 9 we cover numerical tests of these algorithms. We devote Sec. 10 to conclusions. In Appendix A we specify some small families of matrices whose LRA fails unless it involves all entries of $M$. In Appendix B we first comment on some random sketching LRA and then define Abridged SRHT Ultrasparse matrices. In Appendix C we recall the known estimates for the norms of a Gaussian matrix and its pseudo inverse.
@@
1.7. Some background for LRA. In this subsection, we write $|||\cdot|||$ to unify the spectral norm $||\cdot||$ and the Frobenius norm $||\cdot||_{F}$, by following [39, Thm. 9.1]. Hereafter $\mathbb{R}^{p\times q}$ denotes the class of $p\times q$ real matrices. For five positive integers $k,l,m,n,r$ satisfying Eqn. (1.1), the triplet of matrices
@@
as well as the product $XYZ^{*}$ represent a rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ with the error matrix $E:=M-~XYZ^{*}$ (see Fig. 1).
[…]
Figure 1: LRA of $M$.
One can replace the triplet $\{X,Y,Z^{*}\}$ with the pair $\{XY,Z^{*}\}$ of $m\times\ell$ and $\ell\times n$ matrices or $\{X,YZ^{*}\}$ of $m\times k$ and $k\times n$ matrices.
An important 3-factor LRA of $M$ is its $r$-top SVD (see Fig. 2), $M_{r}=U_{r}\Sigma_{r}V_{r}^{*}$ for a diagonal matrix $\Sigma_{r}=\operatorname{diag}(\sigma_{j})_{j=1}^{r}$ of the $r$ largest singular values of $M$ and two matrices $U_{r}$ and $V_{r}$ of the $r$ associated left and right singular vectors, respectively. $M_{r}$ is said to be the $r$-truncation of $M$. It can be obtained from a matrix $M$ by setting to 0 all its singular values but the $r$ top (largest) ones. $r$-top SVD of $M$ is a compact SVD of $M_{r}$.
MM==00++$E$
Figure 2: The figure represents top SVD of a matrix.
@@
(Eckart-Young-Mirsky Theorem, see [29, Thm. 2.4.8].) It holds that
@@
Theorem 1.2.
[29, Cor. 8.6.2]. For a pair of $m\times n$} matrices $M$ and $M+E$ it holds that
[…]
$M^{+}$ denotes the Moore–Penrose pseudo inverse of $M$.
@@ 2 CUR decomposition and CUR LRA
We call $G$ the generator and $U$ the nucleus of CUR decomposition of $M$ (see Fig. 1 for $X=C$, $Y=U$, $Z=R^{*}$, and $E=O$).
@@
The pioneering papers [36, 34, 37], as well as [32, 33, 30, 49], define CGR approximations having nuclei $G$; “G” can stand, say, for “germ”. We use the acronym CUR, which is more customary in the West. “U” can stand, say, for “unification factor”, but notice the alternatives of CNR, CCR, or CSR with $N$, $C$, and $S$ standing for “nucleus”, “core”, and “seed”.
@@
is reached for the nucleus $U=C^{+}MR^{+}$ (see [45, Eqn. (6)]). We, however, cannot compute such a nucleus superfast and instead seek canonical CUR LRA (cf. [20, 16, 49]) whose nucleus is the Moore-Penrose pseudo inverse of the $r$-truncation of a given CUR generator:
@@
2.1 CUR decomposition of a matrix
We do not use the following result, but it can be of independent interest.
@@ 2.1 CUR decomposition of a matrix
Proof.
Now let $M=M^{\prime}=CUR$. Then clearly
Hence
It remains to deduce that
but in this case $G_{r}=G$, and so
Remark 2.2.
$M=\begin{pmatrix}1&0&0\\ 0&\epsilon&0\\ 0&1&0\end{pmatrix}$,
3 Superfast CUR LRA algorithms
Next we specify some superfast CUR LRA algorithms as heuristic but estimate their output accuracy in Thm. 4.1.
@@ 3.1 Primitive and Cynical algorithms
Given five integers $k$, $l$, $m$, $n$, and $r$ satisfying (1.1) and an $m\times n$ matrix $M$, sample uniformly or fix a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $k$ row and $l$ column indexes, respectively, to define a CUR generator $G$, compute its $r$-truncation $G_{r}$, build on it a canonical CUR LRA of $M$, and call this LRA algorithm Primitive. Apart from the selection of the sets $\mathcal{I}$ and $\mathcal{J}$ the algorithm amounts to computing the $r$-truncation $(M_{\mathcal{I},\mathcal{J}})_{r}$ of the matrix $M_{\mathcal{I},\mathcal{J}}$ and its pseudo inverse $((M_{\mathcal{I},\mathcal{J}})_{r})^{+}$; this involves about $kl$ scalars and $O(kl\min\{k,l\})$ flops.
@@
fix or randomly sample a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $p$ row and $q$ column indexes, respectively, compute a $k\times l$ CUR generator $G_{k,l}$ for the $p\times q$ submatrix $M_{\mathcal{I},\mathcal{J}}$ by applying to it one of the algorithms of [26, 50, 20], compute the $r$-truncation $G=G_{k,l,r}$, and build on it CUR LRA of $M$.
@@ 3.2 Cross-Approximation (C–A) iterations
C–A iterations (see Fig. 4):
  • For an $m\times n$ matrix $M$ and target rank $r$, fix four integers $k$, $l$, $p$ and $q$ satisfying (3.1). [C-A iterations are simplified in a special case where $p:=k$ and $q:=l$.]
  • Fix an $m\times q$ “vertical” submatrix of the matrix $M$, made up of its $q$ columns, [^9] One can alternatively begin C–A iterations with a “horizontal” submatrix. and apply to it one of the algorithms of [26, 50, 20], which output a $p\times q$ submatrix $G$.
  • Apply to it the Cynical algorithm and extend its output to CUR LRA of $M$.
  • Swap $p$ and $q$ and reapply the algorithm to the matrix $M^{*}$.
    [This is equivalent to computing a $k\times l$ CUR generator of a fixed $p\times n$ “horizontal” submatrix $M_{1}$ of $M$ that covers the submatrix $G$.]
  • Recursively alternate such “vertical” and “horizontal” steps until the new CUR generator computed at the current iteration coincides with the original one or until the number of recursive C–A steps exceeds a fixed tolerance bound. In both cases output the current CUR LRA of $M$.
Computational Cost. At its every “vertical” (resp. “horizontal”) step the algorithm involves about $mq$ (resp. $nl$) scalars and $O(mq^{2})$ (resp. $O(nl^{2})$) flops. For $p=q=r$ the algorithm of [30] uses $O(mr)$ (resp. $O(nr)$) flops per a “vertical” (resp. “horizontal”) step. Extensive effort has been invested in estimating the required number of C-A iterations (see [32, 30, 49, 1] and the references therein), but adequate support of their superfast empirical performance is still a challenge. [^10] Goreinov et al proved in [30] that every C-A step of maxvol (their variant of C-A iterations in the case where $p=q=r$) monotone increases the volume $|\det(G)|$, and reasonable bounds on the output error norms at a global maximum of the volume have been proved in [32, 49, 1]. Proving that maxvol yields global maximum superfast is still a challenge, however.
Figure 4: The first three recursive C–A steps output three striped submatrices.
Initialization recipes. Adaptive C-A iterations are initialized based on Gaussian elimination with complete pivoting (at a superlinear cost) combined with dynamic search for gaps in the spectrum of the singular values of $M$. Empirically the iterations converge very fast for a large and important class of input coming from the study of ODE and PDE (cf. [2, 3, 5, 6]). Maxvol variant of C-A iterations of [30] for $q=p=k=l=r$ is initialized by using $O(nr^{2})$ flops and still empirically converges fast for a large class of inputs, although the authors suggest using Gaussian elimination with complete pivoting if their initialization recipe does not work. One can try various other heuristic initialization recipes, e.g., based on progressive alternating direction pivoting of the papers [43, 74, 19].
3.3 Maxvol for rank-1 approximation and 1- and $\infty$-norm estimation
@@
In that case Eqns. (3.2) and (3.3) imply that the 1-norms of the $j$th column vector $C:=(m_{g,j})_{g=1}^{m}$ and the $i$th row vector $R:=(m_{i,h})_{h=1}^{n}$ estimate the 1-norm and the $\infty$-norm of $M$, respectively, being off by at most factors of $m$ and $n$, respectively.
@@ 4.1 Error Estimation: an Outline and the Statement
We estimate the error norm of CUR rank-$r$ approximation of $M$ by comparing it with CUR decomposition of a nearby rank-$r$ matrix $M^{\prime}$ defined by the same sets of row and column indices.

### v3

@@ 1 Introduction
1.5. Organization of our paper. We recall some background material on LRA, Gaussain and factor-Gaussian matrices, and CUR LRA in the next two subsections and Sec. 2, respectively. In Part I (Secs. 26) we study computation of CUR LRA. In Sec. 3 we briefly comment on C-A iterations and specify some more rudimentary superfast CUR LRA. In Sec. 4 we estimate output errors of any CUR LRA, thus proving our Main Result 1. In Sec. 5 we a little strengthen it in the case of a perturbed factor-Gaussian input matrix. We devote Sec. 6 to our numerical experiments for the algorithms of Part I. In Part II (Secs. 710) we study random sketching algorithms for LRA. We recall and a little extend some known ones in Sec. 7 and recall their deterministic output error bounds in Sec. 8. In Sec. 9 we prove our Main Results 2 and 3, on dual output error bounds. In Sec. 10 we cover numerical tests of the algorithms of Part II. We devote Sec. 11 to conclusions. In Appendix A we specify some small families of matrices on which any LRA fails miserably unless it involves all entries of $M$. In Appendix B we first comment on some random sketching LRA and then define Abridged SRHT Ultrasparse matrices. In Appendix C we recall the estimates for the norms of a Gaussian matrix and its pseudo inverse. In Appendix D we prove that pre- and post-multiplication by Gaussian matrices turns any rank-$r$ matrix into a factor-Gaussian matrix having expected rank $r$. In Appendix E we prove necessary and sufficient condition for having CUR decomposition of a matrix.
In Secs. 3 and 7 we describe large families of LRA algorithms covered by our Main Results. In Secs. 6 and 10 we tested numerically some specified algorithms from these families, which supported our Main Results empirically. This was the only purpose of our numerical tests, but the observation of potential benefits of using $d$-Abridged SRHT multipliers has come as a by-product of that study.
@@
1.6. Some background for LRA. In this subsection, we write $|||\cdot|||$ to unify the spectral norm $||\cdot||$ and the Frobenius norm $||\cdot||_{F}$, by following [30, Thm. 9.1].
Hereafter $\mathbb{R}^{p\times q}$ denotes the class of $p\times q$ real matrices.
For positive integers $k,l,m,n,r$ satisfying (1.1), the triplet of matrices
@@
as well as the product $XYZ^{*}$ represent a rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ with the error matrix $E:=M-~XYZ^{*}$. One can replace the triplet $\{X,Y,Z^{*}\}$ with the pair $\{XY,Z^{*}\}$ of $m\times\ell$ and $\ell\times n$ matrices or $\{X,YZ^{*}\}$ of $m\times k$ and $k\times n$ matrices.
An important 3-factor LRA of $M$ is its $r$-top SVD $M_{r}=U_{r}\Sigma_{r}V_{r}^{*}$ for a diagonal matrix $\Sigma_{r}=\operatorname{diag}(\sigma_{j})_{j=1}^{r}$ of the $r$ largest singular values of $M$ and two matrices $U_{r}$ and $V_{r}$ of the $r$ associated left and right singular vectors, respectively.
$M_{r}$ is said to be the $r$-truncation of $M$. One can obtain it from a matrix $M$ by setting to 0 all its singular values but the $r$ top (largest) ones. $r$-top SVD of $M$ is compact SVD of $M_{r}$.
$M^{+}$ denotes the Moore–Penrose pseudo inverse of $M$.
@@
(Eckart-Young-Mirsky Theorem, see [22, Thm. 2.4.8].) It holds that
@@
Theorem 1.2.
[The impact of a perturbation of a matrix on its singular values (see [22, Cor. 8.6.2]).] For a pair of $m\times n$} matrices $M$ and $M+E$ it holds that
[…]
Lemma 1.2.
(The impact of a perturbation of a matrix on its singular space, adapted from [63], [55, Thm. 6.4], [26, Thm. 1].) Let $M$ be an $m\times n$ matrix of rank $r<\min(m,n)$ where
$M=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}~\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}~\begin{bmatrix}V_{r}^{T}\\ V_{\perp}^{T}\end{bmatrix}$
is its SVD, and let $E$ be a perturbation matrix such that
[…]
Remark 1.1.
Matrix $\tilde{V}$ above does not necessarily have orthogonal columns, but one can readily prove that the matrix $(V_{r}+V_{\perp}P)(I_{r}+P^{T}P)^{-1/2}$ has orthonormal columns, i.e., $(I_{r}+P^{T}P)^{-1/2}$ normalizes $\tilde{V}$ (see [63, 55, 26]).
1.7. Background on Gaussian and factor-Gaussian matrices. Constant matrices are filled with constants, unlike random matrices, filled with random variables.
Theorem 1.3.
Proof.
Rank deficiency of matrices $F$, $H$, FMFM, and $MH$ is equivalent to turning into 0 the determinants $\det(FF^{*})$, $\det(H^{*}H)$, $\det(FM(FM)^{*})$, and $\det((MH^{*})MH)$, respectively. The claim follows because these equations define algebraic varieties of lower dimension in the linear spaces of the entries, considered independent variables (cf., e.g., [6, Prop. 1]). ∎
Remark 1.2.
Events that occur with probability 0 are immaterial for our probability estimates, and hereafter we say that a matrix has rank $r$ if it is has rank $r$ with probability 1.
Lemma 1.3.
[Orthogonal Invariance.] [58, Theorem 3.2.1]. Suppose that $G$ is an $m\times n$ Gaussian matrix, $k\leq\min\{m,n\}$ is a positive integer, and $S\in\mathbb{R}^{k\times m}$ and $T\in\mathbb{R}^{n\times k}$ are constant matrices, having orthonormal rows and columns, respectively. Then $SG$ and $GT$ are random matrices having distribution of $k\times n$ and $m\times k$ Gaussian random matrices, respectively.
Definition 1.1.
[Factor-Gaussian matrices.] Let $A\in\mathbb{R}^{m\times r}$, $B\in\mathbb{R}^{r\times n}$, and $C\in\mathbb{R}^{r\times r}$ be three constant well-conditioned matrices of full rank $r<\min\{m,n\}$. Let $G_{1}$ and $G_{2}$ be $m\times r$ and $r\times n$ independent Gaussian matrices, respectively. Then $G_{1}B$, $AG_{2}$, and $G_{1}CG_{2}$ are left, right, and two-sided factor-Gaussian matrices of rank $r$, respectively.
Theorem 1.4.
The distribution of two-sided $m\times n$ factor-Gaussian matrices $G_{m,r}CG_{r,n}$ does not change if in its definition we replace the factor $C$ by an appropriate diagonal matrix $\Sigma=(\sigma_{j})_{j=1}^{r}$ such that $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{r}>0$.
Proof.
Let $C=U_{C}\Sigma_{C}V_{C}^{*}$ be SVD. Then $A=G_{m,r}U_{C}$ and $B=V_{C}^{*}G_{r,n}$ have distributions of independent $m\times r$ and $r\times n$ Gaussian matrices, respectively, by virtue of Lemma 1.3. Hence $G_{m,r}CG_{r,n}$ has the same distribution as $A\Sigma_{C}B$. ∎
@@ 2 CUR decomposition and CUR LRA
We call $G$ the generator and $U$ the nucleus of CUR decomposition of $M$ where $X=C$, $Y=U$, $Z=R^{*}$, and $E=O$).
@@
The pioneering papers of 1995-2001 (cf. [27, 25] and the references in [40]) define CGR approximations having nuclei $G$; “G” can stand, say, for “germ”. We use the acronym CUR, which is more customary in the West. “U” can stand, say, for “unification factor”, but notice the alternatives of CNR, CCR, or CSR with $N$, $C$, and $S$ standing for “nucleus”, “core”, and “seed”.
@@
is reached for the nucleus
$U=C^{+}MR^{+}$ (2.2)
(see [36, Eqn. (6)]). We, however, cannot compute such a nucleus superfast and instead seek canonical CUR LRA (cf. [16, 40]) whose nucleus is the Moore-Penrose pseudo inverse of the $r$-truncation of a given CUR generator:
@@
We do not use the following result, but it can be of independent interest. We prove it in Appendix E.
@@
3 Some superfast CUR LRA algorithms
Next we present some superfast CUR LRA algorithms as heuristic. Their crude output accuracy estimates in Thm. 4.1 are greatly superseded in our numerical tests of their specified versions in Sec. 6.
@@ 3.1 Primitive and Cynical algorithms
Given five integers $k$, $l$, $m$, $n$, and $r$ satisfying (1.1) and an $m\times n$ matrix $M$, sample uniformly or fix a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $k$ row and $l$ column indexes, respectively, to define a CUR generator $G$, compute its $r$-truncation $G_{r}$, build on it a canonical CUR LRA of $M$, and call this LRA algorithm Primitive. Given the sets $\mathcal{I}$ and $\mathcal{J}$, the algorithm amounts to computing the $r$-truncation $(M_{\mathcal{I},\mathcal{J}})_{r}$ of the matrix $M_{\mathcal{I},\mathcal{J}}$ and its pseudo inverse $((M_{\mathcal{I},\mathcal{J}})_{r})^{+}$; this involves about $kl$ scalars and $O(kl\min\{k,l\})$ flops.
@@
uniformly sample a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $p$ row and $q$ column indexes, respectively, compute a $k\times l$ CUR generator $G_{k,l}$ for the $p\times q$ submatrix $M_{\mathcal{I},\mathcal{J}}$ by applying to it one of the algorithms of [19, 41, 16], compute the $r$-truncation $G=G_{k,l,r}$, and build on it CUR LRA of $M$.
@@ 3.2 Cross-Approximation (C–A) iterations
We refer to [23, 3, 40, 1], and the references therein for variation and performance of C-A iterations. In particular their maxvol version in [23] involves $O((m+n)r)$ flops per C-A iteration seeking an $r\times r$ CUR generator for an $m\times n$ matrix. Extensive effort has been invested into estimation of the overall number of C-A iterations required for convergence (see [25, 23, 40, 1] and the references therein), but adequate support of their superfast empirical performance is still a challenge. [^10] Goreinov et al proved in [23] that every C-A step of maxvol monotone increases the volume $|\det(G)|$, and reasonable bounds on the output error norms at a global maximum of the volume have been proved in [25, 40, 1]. Proving that maxvol yields global maximum superfast is still a challenge, however, although such bounds hold whp where an input matrix $M$ lies close to a factor-Gaussian matrix (see [47]).
Initialization recipes. Adaptive C-A iterations are initialized based on Gaussian elimination with complete pivoting (at a superlinear cost) combined with dynamic search for gaps in the spectrum of the singular values of $M$. Empirically the iterations converge very fast for a large and important class of inputs coming from the study of ODE and PDE (cf. [3] and the references therein). Maxvol variant of C-A iterations of [23] for $q=p=k=l=r$ is initialized by using $O(nr^{2})$ flops and still empirically converges fast for a large class of inputs, although the authors admit that their initialization recipe may not work and then suggest using costly Gaussian elimination with complete pivoting. For heuristic initialization one can alternatively try using the outputs of the algorithms of [34, 65, 15].
3.3 Maxvol for rank-1 approximation and norm estimation
@@
In that case Eqns. (3.2) and (3.3) imply that the 1-norms of the $j$th column vector $C:=(m_{g,j})_{g=1}^{m}$ and the $i$th row vector $R:=(m_{i,h})_{h=1}^{n}$ provide crude estimates for the 1-norm and the $\infty$-norm of $M$, respectively, being off by at most factors of $m$ and $n$, respectively.
@@ 4.1 Error Estimation: an Outline and the Statement
We estimate the error norm of CUR rank-$r$ approximation of $M$ by comparing it with CUR decomposition of a nearby rank-$r$ matrix $M^{\prime}$ defined by the same sets of row and column indexes.
@@
5 The errors of CUR LRA of a perturbed factor-Gaussian matrix
In this section we prove that a canonical CUR LRA of a matrix $M$ with a generator $G=M_{\mathcal{I},\mathcal{J}}$ is quite accurate whp for any fixed pair of sufficiently large sets $\mathcal{I}$ and $\mathcal{J}$ of rows and columns indexes if $M$ is close to a two-sided factor-Gaussian matrix $M^{\prime}$ of low rank. Unlike Def. 1.1 we use character $H$ for Gaussian matrices, leaving character $G$ for CUR generators.
Theorem 5.1.
Given two independent Gaussian matrices $H_{1}\in\mathbb{R}^{m\times\rho}$ and $H_{2}\in\mathbb{R}^{\rho\times n}$ and a $\rho\times\rho$ (well-conditioned) matrix $\Sigma$ of full rank $\rho$ such that
$k<m,~l<n,~\textrm{and}~\min\{k,l\}\geq\rho+2\geq 4,$ (5.1)
write
$M^{\prime}:=H_{1}\Sigma H_{2},$ (5.2)
$C^{\prime}:=M^{\prime}_{:,\mathcal{J}},~R^{\prime}:=M^{\prime}_{\mathcal{I},:},~G^{\prime}=M^{\prime}_{\mathcal{I},\mathcal{J}}\in\mathbb{R}^{k\times l},~{\rm and}~U^{\prime}:=G^{\prime+}$
for fixed row and column index sets $\mathcal{I}$ and $\mathcal{J}$. Then $C^{\prime}$, $R^{\prime}$, and $G^{\prime}$ are factor-Gaussian matrices such that
[…]
for  $e:=2.7182818\dots$,
$\alpha:=\frac{e^{2}\sqrt{kl}}{(k-\rho)(l-\rho)}~\textrm{and}~\beta:=\max\big{\{}(\sqrt{m}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{l}),(\sqrt{n}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{k})\big{\}}.$ (5.3)
Proof.
[…]
Recall that $C^{\prime}=M^{\prime}_{:,\mathcal{J}}=H_{1}\Sigma{H_{2}}_{:,\mathcal{J}}$ and $H_{1}$ and $H_{2}$ are independent of one another and deduce that
[…]
Recall from Thm. C.1 that $\mathbb{E}(\nu_{p,q})\leq\sqrt{p}+\sqrt{q}$ and obtain
[…]
Next estimate $U^{\prime}||$. Apply Lemma 1.1, recall from Thm. C.2 that
$\mathbb{E}(\nu^{+}_{p,q})\leq\frac{e\sqrt{p}}{p-q},~{\rm for}~p\geq q+2\geq 2,$
and deduce that
[…]
$\leq\frac{e^{2}\sqrt{kl}}{(k-\rho)(l-\rho)\sigma_{\rho}(\Sigma)}=\frac{\alpha}{\sigma_{\rho}(\Sigma)}.$
Now assume that an input matrix $M$ is a perturbed two-sided factor-Gaussian matrix and that two fixed sets $\mathcal{I}$ and $\mathcal{J}$ of its row and column indexes are sufficiently large and then estimate the norms $U||$, $C||$ and $R||$.
Lemma 5.1.
Under the assumptions of Thm. 5.1, let $\epsilon\leq\frac{\sigma_{\rho}(\Sigma)}{30\alpha}$ be a small positive number and let $E\in\mathbb{R}^{m\times n}$ be a perturbation matrix such that $E||\leq\epsilon$. Write
$M:=M^{\prime}+E,~C:=M_{:,\mathcal{J}},~R:=M_{\mathcal{I},;},~{\rm and}~U:=(M_{\mathcal{I},\mathcal{J}})_{\rho}^{+}.$
Then with a probability no less than 0.7 we have
[…]
Proof.
[…]
$U^{\prime}||\leq\frac{10\alpha}{\sigma_{\rho}(\Sigma)}\textrm{ or equivalently }\sigma_{\rho}(M^{\prime}_{\mathcal{I},\mathcal{J}})\geq\frac{\sigma_{\rho}(\Sigma)}{10\alpha}$ (5.4)
with a probability no less than 0.9.
Similarly to the argument of Lemma 4.3, deduce that
$\sigma_{\rho}(M_{\mathcal{I},\mathcal{J}})\geq\sigma_{\rho}(M^{\prime}_{\mathcal{I},\mathcal{J}})-\epsilon\geq\frac{\sigma_{\rho}(\Sigma)}{15\alpha},$
and then claim (i) follows.
[…]
with a probability no less than 0.8. Then combine claim (i) with the union bound and obtain that
[…]
with a probability no less than 0.7. ∎
Now choose any generator of CUR LRA of a perturbed factor-Gaussian matrix of rank $\rho$ and prove a probabilistic upper bound on the error of a resulting CUR LRA of that matrix.
Theorem 5.2.
If the assumptions of Lemma 5.1 hold for a perturbed factor-Gaussian matrix $M$, then bound (4.5) on the approximation error norm MCUR||M-CUR|| for $\theta\leq 1/2$, that is,
$M-CUR||\leq(4\zeta v+2)(v+1)\epsilon,~{\rm for}~4\zeta\leq 2+2\sqrt{5},~v=150\alpha\beta\sigma_{1}(\Sigma)/\sigma_{\rho}(\Sigma)+1/2,$ (5.5)
holds with a probability no less than 0.6.
Proof.
Combine bounds $\epsilon\leq\frac{\sigma_{\rho}(\Sigma)}{30\alpha}$ and (5.4) of Lemma 5.1 and obtain that $\sigma_{\rho}(G)\geq 3\epsilon>\epsilon$ and therefore (4.5) hold with a probability $p_{0}\geq 0.9$. Replace $\theta$ and $v$ with the upper bounds 1/2 and $150\alpha\beta\sigma_{1}(\Sigma)/\sigma_{\rho}(\Sigma)+1/2$, respectively, holding with a probability $p_{1}\geq 0.7$. Combine this bound with $p_{0}\geq 0.9$. ∎
Remark 5.1.
Recall that
$\alpha:=\frac{e^{2}\sqrt{kl}}{(k-\rho)(l-\rho)}~{\rm and}~\beta:=\max\big{\{}(\sqrt{m}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{l}),(\sqrt{n}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{k})\big{\}}.$
Furthermore, suppose that $m\gg k\gg\rho$, $n\gg l\gg\rho$, and
$v=150\alpha\beta\sigma_{1}(\Sigma)/\sigma_{\rho}(\Sigma)+1/2=O\Big{(}\max\Big{\{}\frac{\sqrt{m}}{\sqrt{k}},\frac{\sqrt{n}}{\sqrt{l}}\Big{\}}\Big{)}.$
Now drop the smaller terms of (5.5), and obtain that its dominant part is
$O\Big{(}\max\Big{\{}\frac{m}{k},\frac{n}{l}\Big{\}}\cdot\epsilon\Big{)}.$